# HiddenDetect: Detecting Jailbreak Attacks against Large Vision-Language Models via Monitoring Hidden States

### v2

@@ 5.1.1 Dataset and models
We consider realistic scenarios where both text-based attack and bi-modal attack could happen. For text-based attack evaluation, two datasets are considered. The first, XSTest Röttger et al. (2024), is a test suite containing 250 safe prompts across 10 categories and 200 crafted unsafe prompts. This dataset is widely used to assess the performance of methods against text-based LVLM attacks. The second dataset, FigTXT, was specifically developed for this study. It comprises instruction-based text jailbreak queries extracted from the original FigStep Gong et al. (2023) dataset, serving as malicious user queries. In addition, a corpus of 300 benign user queries was constructed, with further details on its creation provided in the Appendix.
For bi-modal attack, the test set is also constructed to include both unsafe and safe examples. Unsafe examples are sourced from MM-SafetyBench Liu et al. (2023c), a dataset comprising typographical images, stable diffusion-generated images, Typo + SD images, and text-based attack samples. Additional unsafe examples are derived from FigIMG, which includes typographical jailbreak images and paired prompts targeting ten toxic themes from the original FigStep Gong et al. (2023) dataset. Safe examples are drawn from MM-Vet, a benchmark designed to assess core LVLM capabilities, such as recognition, OCR, and language generation. The entire MM-Vet dataset is included in both FigIMG and the overall test set to ensure robust coverage of benign scenarios.
We evaluate our method on three popular LVLMs, including LLaVA-1.6-7B Liu et al. (2023a), CogVLM-chat-v1.1 Wang et al. (2023), and Qwen-VL-Chat Bai et al. (2023).
@@ 5.1.2 Baselines and Evaluation Metric
We evaluate the proposed method against a diverse set of baseline approaches, categorized as follows: (1) Uncertainty-based detection methods, including Perplexity (Alon and Kamfonas, 2023), GradSafe (Xie et al., 2024), and Gradient Cuff (Hu et al., 2024); (2) LLM-based approaches, such as Self Detection (Gou et al., 2024) and GPT-4V (OpenAI, 2023); (3) Mutation-based methods, represented by JailGuard (Zhang et al., 2023); and (4) Denoising-based approaches, including MirrorCheck (Fares et al., 2024) and CIDER (Xu et al., 2024).
@@ 5.2 Main Results
The experimental results in Table 1 demonstrate that the proposed method consistently outperforms existing approaches across multiple multimodal large language models (LVLMs) and benchmarks. For LLaVA, CogVLM, and Qwen-VL, it achieves the highest AUROC scores across all datasets, including XSTEST, FigTxt, FigImg, and MM-SafetyBench. These results highlight the effectiveness of the proposed approach in improving performance across diverse models and evaluation settings. When compared to baseline methods, our approach performs better consistently. Simple methods such as Perplexity and Self-detection have much lower average AUROC scores, between 0.638 and 0.734 across the three LVLMs. Even more advanced methods like GradSafe and Gradient Cuff fall short of our performance. For example, Gradient Cuff achieves average AUROC scores of 0.791, 0.769, and 0.716 on LLaVA, CogVLM, and Qwen-VL, while ours achieves 0.922, 0.903, and 0.826. This shows that our method is much more effective at integrating reasoning across text and image inputs. Our method’s ability to perform well on various VLMs shows that it works well across different architectures without requiring extra modifications, and is practical for improving the safety of LVLMs in a wide range of scenarios.
@@ 5.3 Ablation Study
Effect of the Most Safety-Aware Layers. To assess their role in HiddenDetect, we compare three settings: (1) exclusion of these layers, (2) aggregation across all layers, and (3) the original setting, which focuses on them. Detection performance is measured using AUPRC. Unlike Section 5.1, which employs trapz AUC, this ablation study uses simple summation for fairness, with negligible impact on overall performance. Table 2 shows that the original setting consistently outperforms both variants, especially when excluding these layers.
Effect of Scaling the Weights of Safety-Aware Layers. Using our few-shot approach, we identify layers 16–29 as the Most Safety-Aware Layers in LLaVA-v1.6-Vicuna-7B. To validate their role in safety performance, we adopt the methodology from Li et al. (2024a), which evaluates layer impact by analyzing changes in over-rejection rates for benign queries containing certain malicious words when layer weights are scaled. We extend this analysis by incorporating paired benign images to create a bimodal evaluation dataset (details in the appendix). As shown in Table 3, increasing the scaling factor for these layers results in a higher number of rejected samples, with scaling all layers within this range yielding the highest rejection count for both scaling factors.
@@ 5.4 Visualization
We demonstrate HiddenDetect’s effectiveness by projecting the last token’s hidden state logits onto a plane defined by the Refusal Vector and an orthogonal vector capturing the query’s semantics. We use LLaVA v1.6 Vicuna 7B with bimodal jailbreak samples from Figstep, contrasts toxic (red) and benign (blue) samples from MM-Vet. As shown in Figure 4, early layers exhibit a mixed distribution of red and blue dots along the refusal semantic dimension. By layer 10, toxic samples shift toward the refusal direction, with the greatest separation at layers 22, 23, and 24. In these layers, benign queries exhibit stronger refusal projections. Notably, despite higher projections in the final layer, many malicious queries still show lower refusal scores than benign ones, revealing classification inconsistencies.
@@ A.3 Safe samples in FigTxt
The safe samples in FigTxt consist of half direct safe questions and half safe questions structured similarly to FgStep texts, employing a step-by-step questioning format.

### v4

@@ 5.1.1 Dataset and models
We consider realistic scenarios where both text-based attack and bi-modal attack could happen. For text-based attack evaluation, two datasets are considered. The first, XSTest Röttger et al. (2024) is a test suite comprising 250 safe prompts and 200 crafted unsafe prompts. It is widely used to identify exaggerated safety behaviors in LVLMs. We use it to evaluate how frequently our method incorrectly flags safe prompts. The second dataset, FigTxt, was specifically developed for this study. It comprises instruction-based text jailbreak queries extracted from the original FigStep Gong et al. (2023) dataset, serving as malicious user queries. In addition, a corpus of 300 benign user queries was constructed, with further details on its creation provided in the Appendix A.3. For the bi-modal attack setting, the test set is constructed to include both unsafe and safe examples. Unsafe examples are sourced from MM-SafetyBenchLiu et al. (2023b) and JailBreakV-28KLuo et al. (2024b), which comprise typographical images, Stable Diffusion-generated images, Typo + SD combinations, and text-based attack prompts. Additional unsafe examples are derived from FigImg, which includes typographical jailbreak images and paired prompts targeting ten toxic themes from the original FigStep Gong et al. (2023) dataset. Safe examples are drawn from MM-Vet Yu et al. (2024), a benchmark designed to assess core LVLM capabilities such as recognition, OCR, and language generation. The full MM-Vet dataset is included in both FigImg and the overall test set to ensure robust coverage of benign scenarios. We evaluate our method on three popular LVLMs, including LLaVA-1.6-7B Liu et al. (2023a), CogVLM-chat-v1.1 Wang et al. (2023), and Qwen-VL-Chat Bai et al. (2023).
@@ 5.1.2 Baselines and Evaluation Metric
We evaluate the proposed method against a diverse set of baseline approaches, categorized as follows: (1) Uncertainty-based detection methods, including Perplexity (Alon and Kamfonas, 2023) and GradSafe (Xie et al., 2024); (2) LLM-based approaches, such as Self Detection (Gou et al., 2024) and GPT-4V (OpenAI, 2023); (3) Mutation-based methods, represented by JailGuard (Zhang et al., 2023); and (4) Denoising-based approaches, including MirrorCheck (Fares et al., 2024) and CIDER (Xu et al., 2024).
@@ 5.2 Main Results
The experimental results in Table 1 demonstrate that the proposed method consistently outperforms existing approaches across multiple multimodal large language models (LVLMs) and benchmarks. For LLaVA, CogVLM, and Qwen-VL, it achieves the highest AUROC scores across all datasets, including XSTest, FigTxt, FigImg, and MM-SafetyBench. These results highlight the effectiveness of the proposed approach in improving performance across diverse models and evaluation settings. When compared to baseline methods, our approach performs better consistently. Simple methods such as Perplexity and Self-detection have much lower average AUROC scores, between 0.647 and 0.748 across the three LVLMs. Even more advanced methods like GradSafe and JailGuard fall short of our performance. For example, GradSafe achieves an average AUROC of 0.777 across the three models evaluated on all test sets, whereas our method attains 0.888. This demonstrates that our method is significantly more effective at detecting malicious intent across both text and image modalities. Our method’s ability to perform well on various LVLMs shows that it works well across different architectures without requiring extra modifications, and is practical for improving the safety of LVLMs in a wide range of scenarios.
@@ 5.3 Ablation Study
Effect of the Most Safety-Aware Layers. To assess their role in HiddenDetect, we compare three settings: (1) exclusion of these layers, (2) aggregation across all layers, and (3) the original setting, which focuses on the most safety-aware layers. Detection performance is measured using AUROC. Unlike Section 5.1, which employs trapz AUC, this ablation study uses simple summation for fairness, with negligible impact on overall performance. Table 2 shows that the original setting consistently outperforms both variants, especially when excluding layers with high safety awareness.
Effect of Scaling the Weights of Safety-Aware Layers. Using our few-shot approach, we identify layers 16–29 as the Most Safety-Aware Layers in LLaVA-v1.6-Vicuna-7B. To validate their role in safety performance, we adopt the methodology from Li et al. (2025a), which evaluates layer impact by analyzing changes in over-rejection rates for benign queries containing certain malicious words when layer weights are scaled. We extend this analysis by incorporating paired benign images to create a bimodal evaluation dataset (details in the Appendix A.4). As shown in Table 3, increasing the scaling factor for these layers results in a higher number of rejected samples, with scaling all layers within this range yielding the highest rejection count for both scaling factors.
@@ 5.4 Visualization
We demonstrate HiddenDetect’s effectiveness by projecting the last token’s hidden state logits onto a plane defined by the Refusal Vector and an orthogonal vector capturing the query’s semantics. We use LLaVA v1.6 Vicuna 7B with bimodal jailbreak samples from Figstep, contrasts toxic (red) and benign (blue) samples from MM-Vet. As shown in Figure 5, early layers exhibit a mixed distribution of red and blue dots along the refusal semantic dimension. By layer 10, toxic samples shift toward the refusal direction, with the greatest separation at layers 22, 23, and 24. In these layers, benign queries exhibit stronger refusal projections. Notably, despite larger projections in the final layer, many malicious queries still show lower refusal scores than benign ones, revealing classification inconsistencies.
@@ A.3 Safe samples in FigTxt
The 300 safe samples in FigTxt consist of 150 shots of direct safe questions and 150 shots of safe questions structured similarly to FigStep texts, employing a step-by-step questioning format.
A.4 Image query used in the second ablation study
We use the below image in our second ablation study to form all bimodal query pairs.